using substitution how do i figure out this problem?
2x+y=-15
y-5x=6

Answer by stanbon(75887)

2x+y=-15
y-5x=6
-----
Solve each for "y":
y = -2x-15
y = 5x+6
----
Substitute and solve for "x":
-2x-15 = 5x+6
7x = -21
x = -3
--------
Solve for "y":
y = 5x+6
y = 5*-3+6
y = -9
---------
Cheers,
Stan H.